# Topological and scheme-theoretic properties of the $D$-graded Proj construction

### v1

@@ Abstract.
Finally, building on Mayeux–Riche’s definition of Serre twists, we give a criterion for their invertibility.
@@ Introduction
There is no natural ample family that recovers the $D$-graded Proj in general, because constructing one requires a choice of a basis of $\operatorname{Pic}(X)$, which is not canonical. As ample families must be ${\mathbb{Z}}^{r}$-graded, we expect that we are unable to recover $\operatorname{Proj}^{D}(S)$ at all when $D$ has torsion (this is already the case for affine toric varieties like the rational normal cone).
Additionally, the components of an ample family must be invertible sheaves on $X=\operatorname{Proj}^{D}(S)$. We were able to give a criterion for the invertibility of twisting sheaves ${\mathcal{O}}_{\operatorname{Proj}^{D}(S)}(d)$ in terms of the finite index subgroups $D^{f}$ arising from relevant $f\in S$.
@@
Organization of the article
In Section 1, we quickly recall the definition and some properties of the $D$-graded Proj construction. Section 2 deals with the generalization of the classical Proj description in terms of $D$-prime ideals and the homogeneous Zariski topology. In Section 3 we present several conditions implying non-separatedness of the Proj construction. Finally, in Section 4, we recall the construction of twisting sheaves for multigraded rings in the lines of [May] and prove a condition for invertibility.
Acknowledgements
Again, I would like to thank my (former) supervisor, Alex Küronya, for his tremendous support and insightful feedback. I am also very grateful to Johannes Horn, Andrés Jaramillo Puentes, Kevin Kühn, Jakob Stix, Martin Ulirsch, and Stefano Urbinati for many helpful discussions and valuable comments.
@@ 3. Separation Criteria
    Let $S={\mathbb{C}}[x,y,z]$ be graded by ${\mathbb{Z}}\times{\mathbb{Z}}/2{\mathbb{Z}}$, such that $\deg(x)=(1,0)$, $\deg(y)=(0,1)$ and $\deg(z)=(1,1)$. In this case, ${\rm Gen}^{D}(S)=\{x,z\}$ (cf. Example 1.5). In contrast to (2), monomic relevant elements have length one, so $h=xy$ cannot be split. In particular, $S$ does not contain a weak element and $\operatorname{Proj}^{D}(S)$ is separated.
@@ 4. Sheaves on Multigraded Spectra
4.1. Invertibility and reflexivity of Serre twists
@@
However, we can detect if ${\mathcal{O}}_{X}(d)$ is a line bundle surprisingly simply.
@@
    ${\mathcal{O}}_{X}(d)$ is invertible as an ${\mathcal{O}}_{X}$-module.
@@
$(a)\implies(b):$ Let ${\mathcal{O}}_{X}(d)$ be invertible, i.e. a line bundle. As $\operatorname{Proj}^{D}(S)$ is covered by the set of all $\operatorname{Spec}(S_{(f)})$ for relevant $f\in S$, the claim follows by the definition of a line bundle.
@@
Let $S$ be an integral domain, $d\in D$ and $(f_{i})_{i\in I}$ a system of generators for $S_{+}$. Then ${\mathcal{O}}_{X}(d)$ is free if and only if $d\in\bigcap_{i\in I}D^{f_{i}}$.
@@
$\Rightarrow$’ By Proposition 4.7 we know that $S(d)_{(f_{i})}$ contains a unit $u_{f_{i}}$ in degree $d$, which is equivalent to $d\in\bigcap_{i\in I}D^{f_{i}}$.
$\Leftarrow$’ Conversely, let $d$ be contained in the intersection of all $D^{f_{i}}$. Then there must exist a unit $u_{i}\mid f_{i}$ of degree $d$ for all $i\in I$ and thus, the map $S_{(f_{i})}\to S(d)_{(f_{i})}$ given by $q\mapsto u_{i}q$ is an isomorphism and turns $S(d)_{(f_{i})}$ into a free $S_{(f_{i})}$-module of rank 1. ∎
@@
For (2) we compute that $D^{x}\cap D^{z}={\mathbb{Z}}\cap({\mathbb{Z}}\times{\mathbb{Z}}/2{\mathbb{Z}})=2{\mathbb{Z}}\times\{0\}$, which shows that ${\mathcal{O}}_{X}(d)$ is invertible if and only if $d=k(2,0)$ for $k\in{\mathbb{Z}}$. However, we cannot exhaust $S$ with the invertible sheaf ${\mathcal{L}}_{S}={\mathcal{O}}_{X}(2,0)$, as the section ring does not contain $y$ or $z$. In particular, one can show that the Picard group of $\operatorname{Proj}^{D}(S)$ is generated by ${\mathcal{O}}_{X}(2,0)$, and hence there cannot exist an ample family recovering $S$.

### v2

@@ Abstract.
Finally, building on Mayeux–Riche’s definition of Serre twists, we give a criterion for their freeness.
@@ Introduction
There is no natural ample family that recovers the $D$-graded Proj in general, because constructing one requires a choice of a basis of $\operatorname{Pic}(X)$, which is not canonical. Concretely, this is since $\operatorname{Cl}(X)$ may contain torsion or has rank greater than one.
Finally, we were able to give a criterion for the freeness of twisting sheaves ${\mathcal{O}}_{\operatorname{Proj}^{D}(S)}(d)$ in terms of the finite index subgroups $D^{f}$ arising from relevant $f\in S$.
@@
Organization of the article. In Section 1, we quickly recall the definition and some properties of the $D$-graded Proj construction. Section 2 deals with the generalization of the classical Proj description in terms of $D$-prime ideals and the homogeneous Zariski topology. In Section 3 we present several conditions implying non-separatedness of the Proj construction. Finally, in Section 4, we recall the construction of twisting sheaves for multigraded rings in the lines of [May] and prove a condition for freeness.
Acknowledgements. Again, I would like to thank my (former) supervisor, Alex Küronya, for his tremendous support and insightful feedback. I am also very grateful to Johannes Horn, Andrés Jaramillo Puentes, Kevin Kühn, Jakob Stix, Martin Ulirsch, and Stefano Urbinati for many helpful discussions and valuable comments.
@@ 3. Separation Criteria
    Let $S={\mathbb{C}}[x,y,z]$ be graded by ${\mathbb{Z}}\times{\mathbb{Z}}/2{\mathbb{Z}}$, such that $\deg(x)=(1,0)$, $\deg(y)=(0,1)$ and $\deg(z)=(1,1)$. In this case, ${\rm Gen}^{D}(S)=\{x,z\}$. In contrast to (2), monomic relevant elements have length one, so $h=xy$ cannot be split. In particular, $S$ does not contain a weak element and $\operatorname{Proj}^{D}(S)$ is separated.
@@ 4. Sheaves on Multigraded Spectra
4.1. Freeness and reflexivity of Serre twists
@@
However, we can detect if ${\mathcal{O}}_{X}(d)$ is a line bundle surprisingly simply, at least for toric $X$. Since $X$ is then $\mathbb{Q}$}-factorial by [paper4], Corollary 2.33. In particular, all ${\mathcal{O}}_{X}(d)$ are invertible.
@@
    ${\mathcal{O}}_{X}(d)$ is free as an ${\mathcal{O}}_{X}$-module.
@@
$(a)\implies(b):$ Let ${\mathcal{O}}_{X}(d)$ be free. As $\operatorname{Proj}^{D}(S)$ is covered by the set of all $\operatorname{Spec}(S_{(f)})$ for relevant $f\in S$, the claim follows by the definition of a line bundle.
@@
Let $S$ be an integral factorially graded domain, $d\in D$ and $(f_{i})_{i\in I}$ a system of generators for $S_{+}$. Then ${\mathcal{O}}_{X}(d)$ is free if and only if $d\in\bigcap_{i\in I}D^{f_{i}}$.
@@
$\Rightarrow$’ By Proposition 4.7 and factoriality, we know that $S(d)_{(f_{i})}$ contains a unit $u_{f_{i}}$ in degree $d$, which is equivalent to $d\in\bigcap_{i\in I}D^{f_{i}}$.
$\Leftarrow$’ Conversely, let $d$ be contained in the intersection of all $D^{f_{i}}$. Then there must exist a unit $u_{i}\in S_{f_{i}}$ of degree $d$ for all $i\in I$ and thus, the map $S_{(f_{i})}\to S(d)_{(f_{i})}$ given by $q\mapsto u_{i}q$ is an isomorphism and turns $S(d)_{(f_{i})}$ into a free $S_{(f_{i})}$-module of rank 1. ∎
@@
For (2) we compute that $D^{x}\cap D^{z}={\mathbb{Z}}\cap({\mathbb{Z}}\times{\mathbb{Z}}/2{\mathbb{Z}})=2{\mathbb{Z}}\times\{0\}$, which shows that ${\mathcal{O}}_{X}(d)$ is free if and only if $d=k(2,0)$ for $k\in{\mathbb{Z}}$. However, as ${\mathcal{O}}_{X}(0,1)^{\otimes 2}={\mathcal{O}}_{X}(0,0)={\mathcal{O}}_{X}$, it is also invertible. Hence, ${\mathcal{O}}_{X}(1,1)$ and ${\mathcal{O}}_{X}(1,0)$ are also invertible. In particular, choosing $O_{X}(1,0)$ and ${\mathcal{O}}_{X}(1,1)$ as members of the ample family, we recover the relevant locus $S_{+}=(x,z)$ and hence its Proj. Note, again, that this choice is not canonical.